# Camouflaged Object Tracking: A Benchmark

### v3

@@ IV Method
In our evaluation, we observed that while HIPTrack [81] outperforms other advanced trackers in COTD, its performance still falls significantly short of satisfactory levels. Camouflaged objects often share similar textures, colors, and patterns with their environment, creating a high level of visual ambiguity. This makes them challenging to detect using single-level feature representation. To overcome this, we propose a novel baseline tracker based on HIPTrack, called HIPTrack-MLS, which enable HIPTrack to sense multi-level features to tackle the difficulties posed by camouflaged objects. By capturing information at various scales and resolutions, multi-level features enhance HIPTrack’s ability to distinguish targets from their environment. This adaptation significantly improves the tracker’s representation power, increasing the chances of accurately identifying camouflaged objects.
@@ IV-A Overall Architecture
Similar to HIPTrack [81], HIPTrack-MLS comprises three main components, as illustrated in Fig. 7 and labeled as (a), (b), and (c), respectively. The feature extraction network is responsible for extracting features from the search region containing target-template matching information while filtering out background image patches. The historical prompt network includes a historical prompt encoder and a historical prompt decoder. The historical prompt encoder encodes the position information and visual features of the target in the current frame as historical target features and stores them in the memory bank of the historical prompt decoder. The historical prompt decoder generates historical prompts and combines them with the compressed features of the search region. The head network has the same structure as OSTrack [95]. The key difference between HIPTrack-MLS and HIPTrack lies in their feature extraction networks. While HIPTrack uses a Vision Transformer (ViT) with early candidate elimination modules, similar to the feature encoder used in OSTrack, HIPTrack-MLS is characterized by a novel Multi-Level Feature Sensing Encoder architecture, as illustrated in Fig. 8. It is designed to enhance feature representation capabilities for camouflaged objects by perceiving features across multiple levels of abstraction. The following section will provide a detailed description of this proposed encoder. Details of other architectures inherited from HIPTrack can be found in [81].
@@ IV-B Multi-Level Feature Sensing Encoder
Our Multi-Level Feature Sensing Encoder is an extension of the Transformer Encoder architecture used in HIPTrack, specifically a Vision Transformer (ViT) equipped with early candidate elimination modules. This structure is comprised of stacked CEBlocks, as illustrated in Fig. 8. The main distinction is the introduction of a Multi-Level Feature Sensing structure for processing multi-level features output by the ViT blocks. In this updated design, features from all levels are directly combined through concatenation, allowing the model to use both detailed information from lower levels and abstract insights from higher levels. A multilayer perceptron efficiently reduces the dimensionality of these multi-level features, using three linear layers with ReLU activation functions to compress and refine the feature representation. Through a process of reshaping and permutation, the module adapts the data to the linear layers, sequentially reducing the dimensionality from 12 to 6, then to 3, and finally to 1. This series of reductions is supplemented by ReLU activation, introducing non-linearity and enhancing the model’s ability to capture complex feature patterns. Given that the early candidate elimination module in each CEBlock might alter the input feature’s shape, we split the CEBlock into two parts: one representing the ViT block with early candidate elimination ($\mathbf{CBlock}^{\rightarrow}_{i}$) and the other without it ($\mathbf{CBlock}^{\downarrow}_{i}$). The corresponding output features are denoted by $H^{\rightarrow}_{i}$ and $H^{\downarrow}_{i}$, respectively. The effect of our proposed Multi-Level Feature Sensing Encoder can be expressed as:
@@ V Evaluation
In this study, precision and success rate serve as the quantitative measures for evaluation.
Precision plot: The precision plot is a standard metric used for evaluating tracking performance. It assesses tracking accuracy by measuring the center position error, which calculates the average Euclidean distance between the tracked object’s center and its ground truth position. This metric summarizes the tracking performance by averaging the errors across all frames in a sequence. However, when a tracker loses the object, its output may become arbitrary, reducing the reliability of the average error [96]. To mitigate this issue, the precision plot is commonly used. It evaluates the overall tracking accuracy by comparing the tracker’s estimated position to the ground truth within a specified distance threshold [96, 97]. In our evaluation, we use a 20-pixel threshold for the precision score, in line with established practices [96].
Success plot: Another evaluation metric commonly used is the success plot. This metric gauges tracker performance by analyzing the overlap between predicted and ground truth bounding boxes. Specifically, for a given tracking bounding box $r_{t}$ and a ground truth bounding box $r_{a}$ , the overlap score is computed as $S=\frac{\left|r_{t}\bigcap r_{a}\right|}{\left|r_{t}\bigcup r_{a}\right|}$ , where $\cap$ and $\cup$ denote the intersection and union of the two bounding boxes, respectively, and ||$\cdot$|| ⋅ | represents the number of pixels contained within these regions. To assess tracking performance across a sequence of frames, we count the number of frames where the overlap score $S$ exceeds a predefined threshold $t_{o}$. The success plot then illustrates the proportion of frames deemed successful at various overlap thresholds ranging from 0 to 1. This visual representation helps in understanding how often the tracker achieves sufficient overlap with the ground truth over different levels of thresholding.Relying on a single success rate at a specific threshold (e.g. $t_{o}$=0.5) may not offer a comprehensive view of a tracker’s performance, as it provides a limited perspective on tracking accuracy. To address this, the area under the curve (AUC) is typically employed in conjunction with the success plot to rank tracking algorithms [96, 97]. The AUC aggregates the success rate across all threshold values, offering a thorough evaluation of the tracking algorithm’s performance across various conditions. This method facilitates a more nuanced and accurate comparison of different tracking algorithms by providing a holistic view of their effectiveness in various scenarios. Each tracker is evaluated using the default parameters provided in the source code. We employ one-pass evaluation (OPE) and assess precision and success rate, as per [98, 99].
V-A Trackers for Comparison
We evaluate our HIPTrack-MLS and 20 state-of-the-art trackers on COTD, including HIPTrack [81], ROMTrack [82], SeqTrack [85], DropTrack [83], AQATrack [100], SparseTT [101], DCPT [102], GRM [86], Stark [103], ZoomTrack [104], MAT [105], SimTrack [87], ARTrack [84], CSWinTT [106], JointNLT [107], TCTrack [108], SiamGAT [109], UDAT [110], HiFT [111], and Aba-ViTrack [112].
V-B Evaluation Results
Overall performance: A comprehensive evaluation was conducted on 20 state-of-the-art trackers and our HIPTrack-MLS using the COTD dataset. The evaluation results are shown in Fig. 9. Notably, HIPTrack-MLS achieves superior performance compared to all other trackers, with a precision (PRC) of 0.525, which is 1.0% higher than its baseline, HIPTrack, which scored 0.515. In terms of success rate (AUC), HIPTrack-MLS also leads with an AUC of 0.692, surpassing the second-place HIPTrack, which has an AUC of 0.680, by 1.2%. These results highlight the outstanding performance of HIPTrack-MLS in both precision and success rate metrics, showcasing its superiority in tackling the challenges of tracking camouflaged objects. This success is largely due to HIPTrack-MLS’s advanced multi-level feature Sensing, which effectively leverages multi-level features.
Figure 11: Qualitative evaluation on 4 sequences from COTD, i.e. fish-4, mantis-2, person-9, and phasmatodea-10 from top to bottom. The results of different methods have been shown with different colors, and ‘GT’ denotes the ground truth.
@@ V-B Evaluation Results
Attribute-based performance: We conducte a comprehensive performance evaluation on eleven common attributes to better analyze and understand the performance of different trackers on the COTD dataset. HIPTrack-MLS consistently ranks first across most attribute subsets. Due to paper length constraints, we present the PRC results for four common challenges in Fig. 10, including Aspect Ration Change (ARC), Illumination Variation (IV), Out-of-View (OV), and Partial Occlusion (POC). As shown, HIPTrack-MLS consistently ranks first across all these attribute subsets, surpassing the baseline HIPTrack by at least 1.0% in all these subsets. Notably, HIPTrack-MLS achieves a precision of 0.271 in the IV subset, surpassing the second place HIPTrack by 3.4%. For the ARC and OV attributes, HIPTrack-MLS shows improvements of 1.3% and 1.0%, respectively, over HIPTrack.
Evaluation on different object categories: In our research, we meticulously conduct a comprehensive experiment to evaluate the performance of multiple state-of-the-art trackers when handling camouflaged objects, with the results systematically presented in Table IV. We strategically categorize the COTD dataset into two superclasses—Aquatic and Non-aquatic—to thoroughly test the capabilities of 20 state-of-the-art trackers, alongside our tracker, HIPTrack-MLS. The experiment hinges on three pivotal metrics: AUC, Precision, and Normalized Precision, to provide a robust assessment of each tracker’s performance. The experimental results demonstrate that our improved tracker HIPTrack-MLS excels in all evaluated metrics, significantly outperforms its baseline, HIPTrack. Specifically, in the domain of non-aquatic object tracking, HIPTrack-MLS achieves an AUC that is 0.58% higher than HIPTrack, with an increase from 63.95% to 64.53%. Additionally, HIPTrack-MLS also leads with a precision of 44.52%, surpassing the second-place HIPTrack, which has a precision of 43.69%, by 0.83%. While there is a minor 0.12% reduction in normalized precision, from 74.22% to 74.10%, this decrement is negligible in the context of real-world applications. When it comes to the aquatic objects, HIPTrack-MLS with a AUC of 77.21%, which is 1.79% higher than its baseline, HIPTrack, which scores 75.42%. For the precision and normalized precision, HIPTrack-MLS shows improvements of 1.08% and 1.48%, respectively, over HIPTrack. These results indicate that our tracker performs exceptionally well in the tracking of camouflaged objects. Additionally, we observe performance variances among other trackers in both objects. Some trackers, such as AQATrack and HIPTrack, show higher precision and normalized precision in aquatic objects, while others, like HiFT and SiamGAT, exhibit relatively lower AUC in non-aquatic objects. This observation potentially indicates limitations in their generalization capabilities and underscores the importance of further research in this area.
TABLE IV: Overall tracking results of the trackers on Non-aquatic and Aquatic,The first-, second- and third-place trackers are labeled with red, blue and green colors respectively.
Tracker Venue Non-aquatic Aquatic
AUC Pre nPre AUC Pre nPre
HiFT [111] ICCV 2021 49.95 29.38 56.19 54.84 25.69 66.90
STARK [103] ICCV 2021 57.93 35.01 66.42 69.80 42.16 84.58
SiamGAT [109] CVPR 2021 50.73 24.45 59.11 59.02 21.79 74.60
SimTrack [87] ECCV 2022 60.02 36.44 67.82 71.01 46.22 85.36
SparseTT [101] IJCAI 2022 57.87 36.00 68.17 69.13 40.93 82.66
CSWinTT [106] CVPR 2022 56.67 32.04 64.04 68.56 43.10 81.76
TCTrack [108] CVPR 2022 38.52 22.68 45.85 46.37 20.87 59.28
UDAT [110] CVPR 2022 49.15 29.61 58.10 59.73 32.75 73.93
Aba-ViTrack [112] ICCV 2023 56.34 33.48 64.80 67.91 40.04 81.56
ROMTrack [82] ICCV 2023 61.16 40.59 71.77 74.90 58.47 87.46
ARTrack [84] CVPR 2023 61.18 40.43 69.80 75.76 58.42 88.83
DropTrack [83] CVPR 2023 63.03 42.61 73.08 75.03 56.14 87.65
GRM [86] CVPR 2023 59.59 38.29 69.06 73.29 53.99 86.56
JointNLT [107] CVPR 2023 44.41 19.94 50.46 58.99 27.45 70.95
MAT [105] CVPR 2023 56.86 33.47 64.35 64.92 37.25 78.03
SeqTrack [85] CVPR 2023 59.23 40.63 70.22 72.29 53.03 86.24
ZoomTrack [104] NeurIPS 2023 58.72 36.86 67.88 72.58 53.72 85.48
DCPT [102] ICRA 2024 60.27 38.71 68.32 73.41 52.04 86.75
AQATrack [100] CVPR 2024 61.74 41.73 71.53 75.72 56.22 88.82
HIPTrack [81] CVPR 2024 63.95 43.69 74.22 75.42 56.42 88.01
HIPTrack-MLS Ours 64.53 44.52 74.10 77.21 57.50 89.49
TABLE V: Efficiency comparison evaluated on COTD.
HIPTrack-MLS HIPTrack ProContEXT ZoomTrack DCPT
FPS 15.67 23.46 10.87 15.62 11.25
TABLE VI: Performance comparison of HIPTrack-MLS and its baseline HIPTrack on generic and our benchmarks.
Tracker LaSOT [98] TrackingNet [99] COTD
Prec. Succ. Prec. Succ. Prec. Succ.
HIPTrack [55] 0.795 0.727 0.838 0.845 0.515 0.680
HIPTrack-MLS 0.793 0.726 0.837 0.845 0.525 0.692
Efficiency analysis: To evaluate the computational efficiency of our proposed method, we conducted runtime analysis on the COTD benchmark dataset. Table Table V presents the tracking speed in frames per second (FPS) for HIPTrack-MLS and other state-of-the-art trackers. The results show that HIPTrack-MLS achieves a tracking speed of 15.67 FPS, which is comparable to ZoomTrack (15.62 FPS) and faster than ProContEXT (10.87 FPS) and DCPT (11.25 FPS). While the original HIPTrack demonstrates higher efficiency at 23.46 FPS, our multi-level feature sensing architecture introduces additional computational overhead due to the feature fusion and dynamic weighting mechanisms. However, this trade-off between speed and accuracy is justified by the significant performance improvements in tracking camouflaged objects, as demonstrated in our previous experiments.
Performance on generic datasets: To validate whether HIPTrack-MLS’s improvements are specifically beneficial for camouflaged object tracking or represent a general enhancement in tracking capability, we conduct comprehensive evaluations on standard tracking benchmarks including LaSOT [98] and TrackingNet [99], alongside our COTD dataset. The results are presented in Table VI. On the LaSOT dataset, HIPTrack-MLS achieves comparable performance to the baseline HIPTrack, with precision and success scores of 0.793 and 0.726, respectively, showing a slight marginal difference compared to HIPTrack’s 0.795 and 0.727. Similarly, on TrackingNet, HIPTrack-MLS demonstrates minimal improvement with precision and success scores of 0.837 and 0.845, compared to HIPTrack’s 0.838 and 0.845. However, on the COTD dataset, HIPTrack-MLS shows a significant performance gain, improving the success rate from 0.680 to 0.692 and precision from 0.515 to 0.525. This notable improvement on COTD, coupled with maintained performance on general tracking benchmarks, suggests that our multi-level feature sensing architecture specifically addresses the challenges of camouflaged object tracking without compromising general tracking capability.
The experimental results support our hypothesis that the multi-level feature sensing mechanism is particularly effective for camouflaged objects, where fine-grained feature discrimination at different scales is crucial for maintaining tracking accuracy. The maintained performance on general tracking benchmarks also indicates that our modifications do not introduce any degradation in general tracking scenarios, making HIPTrack-MLS a versatile tracking solution.
@@
Qualitative evaluation: In Fig. 11, we demonstrate qualitative comparisons of our method with 9 top trackers, including HIPTrack [81], DropTrack [83], DCPT [102], AQATrack [100], ARTrack [84], GRM [86], ROMTrack [82], SeqTrack [85], and ZoomTrack [104]. These examples face challenges such as Illumination Variation, Deformation, Rotation, Aspect Ratio Change, and Partial Occlusion. Specifically, in the fish-4 sequence, only HIPTrack-MLS consistently tracks the camouflaged fish. In mantis-2, while some trackers manage to track the mantis, HIPTrack-MLS consistently demonstrates superior visual performance. In person-9, due to background clutter and multiple object interference, only HIPTrack-MLS, DropTrack, ARTrack, and SeqTrack can track the target person. In phasmatodea-10, many trackers are affected by the background, leading to inaccurate tracking of the camouflaged phasmatodea. These qualitative results validate the superiority of our method for tracking camouflaged objects.
@@
TABLE VII: Illustrates the impact of the importance coefficient of aggregated feature on HIPTrack-MLS’s PRC (%) and AUC (%) on COTD. Red, blue, and green represent the first, second, and third place, respectively.
@@
Figure 12: For each group, the top shows the original images, while the following two rows display the attention map generated by the algorithm in some video frames. The middle row is from HIPTrack, and the bottom row is from HIPTrack-MLS.
V-C Ablation study
Impact of the importance coefficient in Multi-Level Feature Sensing. To examine the effect of the weight coefficient $\gamma$ on balancing the contributions of the aggregated multi-level features and $H_{n}^{\rightarrow}$ in HIPTrack-MLS, we evaluated HIPTrack-MLS on the COTD dataset with a range of $\gamma$ settings. Specifically, $\gamma$ ranges from 0.1 to 1.0 with increments of 0.1. As $\gamma$ increases, the influence of the aggregated feature becomes more pronounced, while the reliance on $H_{n}^{\rightarrow}$ decreases. Table VII details HIPTrack-MLS’s PRC and AUC on the COTD dataset for various $\gamma$ values. The results indicate that the optimal PRC and AUC, which are 52.5% and 69.2% respectively, are achieved when $\gamma$ is set to 0.1. This result indicates that an appropriate choice of $\gamma$ enhances tracking performance, confirming the effectiveness of the proposed method. Consequently, we adopted $\gamma=0.1$ as the default setting for HIPTrack-MLS.
Attention mechanism analysis. To make sense of the advantage of the propose method, we explore the allocation of attention in the task of Camouflaged Object Tracking and compare the performance of the improved HIPTrack-MLS model with the original HIPTrack model. Fig. 12 intuitively displays three types of visual representations: the original image, the attention map of the original HIPTrack model, and the attention map of the HIPTrack-MLS model. The original images in the first row demonstrate the camouflage of the object in a complex background, which poses a significant challenge to tracking algorithms. The high degree of visual integration between the object and the background increases the difficulty of tracking. The second row shows the attention maps generated by the original HIPTrack model. These maps reveal some limitations of the original model when handling camouflaged objects, especially when the similarity between the object and the background is high, the original model may struggle to maintain precise tracking of the object. The third row, however, showcases the attention maps of the HIPTrack-MLS model, which clearly highlight the target object. Even when the object is camouflaged, the model can accurately identify and focus attention on the key features of the object. This improved attention mechanism significantly enhances tracking accuracy, especially in complex environments.
@@ VI Conclusion
In this work, we delve into the relatively underexplored task of tracking camouflaged objects, a field with significant potential for diverse applications. However, there is currently no publicly available benchmark specifically designed for camouflaged object tracking. To address this gap, we introduce the COTD, which is, to the best of our knowledge, the first benchmark specifically designed for camouflaged object tracking. In addition, to evaluate the performance of existing trackers and establish a benchmark for future research, we conducted a comprehensive evaluation and detailed analysis of 20 state-of-the-art tracking algorithms. Additionally, we introduced a new tracker, HIPTrack-MLS, which enhances HIPTrack by improving its multi-level feature sensing capability for better feature representation. Experimental results demonstrate that HIPTrack-MLS significantly outperforms the current leading algorithms on COTD.

### v4

@@ IV Method
In our evaluation, HIPTrack [62] outperforms other advanced trackers in COTD but still falls short of satisfactory levels. Camouflaged objects’ similar textures, colors, and patterns to their environment create visual ambiguity, making single-level feature representation insufficient for detection. To address this, we propose HIPTrack-MLS, a novel baseline tracker based on HIPTrack that enables multi-level feature sensing. By capturing information at various scales and resolutions, multi-level features enhance HIPTrack’s ability to distinguish targets from their environment, significantly improving its representation power and increasing the accuracy of identifying camouflaged objects.
@@ IV-A Overall Architecture
Similar to HIPTrack [62], HIPTrack-MLS comprises three main components, as illustrated in Fig. 5, left panel, and labeled as (a), (b), and (c), respectively. The feature extraction network extracts features from the search region, filtering out background patches. The historical prompt network, with an encoder and decoder, encodes and stores target features and generates prompts combined with search region features. The head network is the same as OSTrack’s. The key difference is in the feature extraction network: HIPTrack-MLS uses a novel Multi-Level Feature Sensing Encoder ( Fig. 5, right panel) to enhance feature representation for camouflaged objects by perceiving multi-level features. Other inherited architectures are detailed in [62].
@@ IV-B Multi-Level Feature Sensing Encoder
Our Multi-Level Feature Sensing Encoder extends the Transformer Encoder architecture in HIPTrack, specifically being a Vision Transformer (ViT) with early candidate elimination modules. It consists of stacked CEBlocks, as shown in Fig. 5, right panel. The key innovation is the Multi-Level Feature Sensing structure that processes multi-level features from ViT blocks. In this design, features from all levels are concatenated, enabling the model to leverage both detailed low-level and abstract high-level information. A multilayer perceptron with three linear layers and ReLU activations reduces the dimensionality from 12 to 6, then to 3, and finally to 1, enhancing the model’s ability to capture complex patterns. To handle potential shape changes due to early candidate elimination, CEBlock is split into two parts: $\mathbf{CBlock}^{\rightarrow}_{i}$ with elimination and $\mathbf{CBlock}^{\downarrow}_{i}$ without it, with corresponding outputs $H^{\rightarrow}_{i}$ and $H^{\downarrow}_{i}$. The effect of our proposed Multi-Level Feature Sensing Encoder can be expressed as:
@@ V Evaluation
We use one-pass evaluation (OPE) to measure precision and success rate, following [77, 78]. Precision is the pixel distance between estimated and ground truth bounding box centers, and success rate is based on IoU exceeding 0.5 [68, 77, 78]. Trackers are ranked by precision at 20 pixels (PRC) and success plot’s AUC.
V-A Evaluation Results
Overall Performance:We evaluated 20 state-of-the-art trackers and our HIPTrack-MLS using the COTD dataset, with results shown in Fig. 6. HIPTrack-MLS outperforms all other trackers, achieving a precision (PRC) of 0.525, 1.0% higher than its baseline HIPTrack (0.515), and an AUC of 0.692, surpassing HIPTrack’s 0.680 by 1.2%. This highlights HIPTrack-MLS’s superior performance in both metrics, thanks to its advanced multi-level feature sensing.
@@ V-A Evaluation Results
Attribute-based performance: We evaluated the performance of different trackers on the COTD dataset across eleven common attributes. HIPTrack-MLS consistently ranks first in most attribute subsets. Due to space constraints, we present PRC results for four challenges in Fig. 7: Aspect Ratio Change (ARC), Illumination Variation (IV), Out-of-View (OV), and Partial Occlusion (POC). More results are in the supplementary material. HIPTrack-MLS ranks first in all these subsets, surpassing HIPTrack by at least 1.0%. Notably, it achieves a precision of 0.271 in IV, beating HIPTrack by 3.4%, and shows improvements of 1.3% in ARC and 1.0% in OV. These results highlight the effectiveness of our method for tracking camouflaged objects.
Figure 8: Qualitative evaluation on 3 COTD sequences: mantis-2, person-9, and phasmatodea-10. Different colors show results of various methods, with ‘GT’ for ground truth.
@@
Qualitative Evaluation: In Fig. 8, we compare our method with 4 top trackers facing challenges like Illumination Variation, Deformation, Rotation, Aspect Ratio Change, and Partial Occlusion. In mantis-2, HIPTrack-MLS shows superior performance. In person-9, only HIPTrack-MLS, HIPTrack, and ARTrack can track the target amidst clutter and interference. In phasmatodea-10, many trackers fail due to background effects, while HIPTrack-MLS remains accurate. These results highlight our method’s effectiveness for tracking camouflaged objects.
V-B Ablation Study
Impact of the importance coefficient in Multi-Level Feature Sensing. We evaluated HIPTrack-MLS on the COTD dataset with $\gamma$ ranging from 0.1 to 1.0 in increments of 0.1 to study its effect on balancing aggregated multi-level features and $H_{n}^{\rightarrow}$. As $\gamma$ increases, the aggregated feature’s influence grows while reliance on $H_{n}^{\rightarrow}$ diminishes. Table III shows PRC and AUC results for different $\gamma$ values. The optimal PRC (52.5%) and AUC (69.2%) are achieved at $\gamma=0.1$, indicating that an appropriate $\gamma$ choice enhances tracking performance. Thus, we set $\gamma=0.1$ as the default for HIPTrack-MLS.
@@ V-B Ablation Study
TABLE III: Shows the effect of the aggregated feature’s importance coefficient on HIPTrack-MLS’s PRC (%) and AUC (%) on COTD. Red, blue, and green denote 1st, 2nd, and 3rd place.
@@ VI Conclusion
This work explores the underexplored task of tracking camouflaged objects, which has significant application potential. Currently, no public benchmark exists for this task. We introduce COTD, the first benchmark for camouflaged object tracking. We also evaluated 20 state-of-the-art tracking algorithms and introduced HIPTrack-MLS, an improved version of HIPTrack with enhanced multi-level feature sensing. Experiments show HIPTrack-MLS outperforms leading algorithms on COTD.